# MI-HET-855

### v1

@@ Abstract
Existing solutions to this anomaly mostly involve additional species in the neutrino sector, utilizing their oscillation, decays, and/or upscattering  Sorel:2003hf ; Karagiorgi:2009nb ; Collin:2016aqd ; Giunti:2011gz ; Giunti:2011cp ; Gariazzo:2017fdh ; Boser:2019rta ; Kopp:2011qd ; Kopp:2013vaa ; Dentler:2018sju ; Abazajian:2012ys ; Conrad:2012qt ; Diaz:2019fwt ; Asaadi:2017bhx ; Karagiorgi:2012kw ; Pas:2005rb ; Doring:2018cob ; Kostelecky:2003cr ; Katori:2006mz ; Diaz:2010ft ; Diaz:2011ia ; Gninenko:2009ks ; Gninenko:2009yf ; Bai:2015ztj ; Moss:2017pur ; Bertuzzo:2018itn ; Ballett:2018ynz ; Fischer:2019fbw ; Moulai:2019gpi ; Dentler:2019dhz ; deGouvea:2019qre ; Datta:2020auq ; Dutta:2020scq ; Abdallah:2020biq ; Abdullahi:2020nyr ; Liao:2016reh ; Carena:2017qhd ; Abdallah:2020vgg . Recently, long-lived mediators and dark matter upscattering have been considered to solve the MiniBooNE puzzle Dutta:2021cip ; PhysRevD.109.095017 . MicroBooNE’s strong capability to distinguish photons from electrons MicroBooNE:2025khi has already constrained the parameter space of some of these solutions, in particular, through a potential signature in the inclusive single-photon channel MicroBooNE:2025ntu but without any obvious excess in the coherent-photon production channel similar to SM MicroBooNE:2025rsd ; Hagaman:2025mrx . Already, these hints are beginning to inform a better understanding of the anomaly as it takes shape.
@@
Both neutrinos and dark matter can be produced from the two-body and three-body decays of charged mesons Dutta:2021cip , respectively, thereby correlating their fluxes to the excess observed in MiniBooNE’s target-mode data through magnetic focusing in the Booster Neutrino Beam (BNB) horn system. This mechanism simultaneously accounts for the absence of an excess in MiniBooNE’s dump-mode observations MiniBooNEDM:2018cxm , where the fluxes are isotropically suppressed due to the lack of magnetic focusing horns Dutta:2021cip .
@@
Model. We consider a simplified model in which a massive scalar couples to quarks, while a massive vector boson couples to either neutrinos, dark matter, or both simultaneously. Under this scenario, the scattering process delineated in Fig. 1 is expected to result in a single photon. A similar process occurs in the SM with neutrinos, via the $Z$-$h$-$\gamma$ interaction. However, the associated cross section and kinematics from these heavy mediators (i.e., $Z$ and $h$) are unsuitable to explain the MiniBooNE excess. This motivates us to introduce a new gauge boson $Z^{\prime}$ and a scalar $\phi$ lighter than the $Z$ and $h$ [^2] This setup can be motivated by the spontaneous breaking of an additional $U(1)$ gauge boson. The effective Lagrangian for $\phi$ and $Z^{\prime}$ with masses $m_{\phi}$ and $m_{Z^{\prime}}$, respectively, is
$\mathcal{L}_{\nu}\supset g_{\nu}\bar{\nu}_{\ell}\gamma^{\alpha}\nu_{\ell}Z_{\alpha}^{\prime}-\frac{g_{\phi Z^{\prime}\gamma}}{2}\phi F_{\mu\nu}{Z^{\prime}}^{\mu\nu}+y_{q}\sum_{q}x_{q}\bar{q}q\phi,$ (1)
@@
where $\ell=e,\mu$, $F_{\mu\nu}$ ($Z^{\prime}_{\mu\nu}$) symbolizes the energy-stress tensor of the SM photon ($Z^{\prime}$) field, $g_{i}$ denotes the associated coupling, $y_{q}$ is the quark flavor-universal coupling, and $x_{q}$ parameterizes the flavor-nonuniveral Yukawa fractions. This Lagrangian can give rise to the scattering process shown in Fig. 1. For the light dark matter solution, we use an alternative (but similar) Lagrangian that generates $\chi\mathcal{N}\to\chi\mathcal{N}\gamma$ for a fermion $\chi$ with mass $m_{\chi}$:
$\mathcal{L}_{\chi}\supset g_{\chi}\bar{\chi}\gamma^{\mu}\chi Z^{\prime}_{\mu}-\frac{g_{\phi Z^{\prime}\gamma}}{2}\phi F_{\mu\nu}{Z^{\prime}}^{\mu\nu}+y_{q}\sum_{q}x_{q}\bar{q}q\phi.$ (2)
@@
In principle, $Z^{\prime}$ can couple to any SM fermions, but for simplicity and illustration, we assume a “neutrinophilic” $Z^{\prime}$; dark matter can be predominantly produced from exotic three-body decays of charged mesons, followed by the $Z^{\prime}$ decay into a $\chi$ pair. Based on the above Lagrangians, we calculate the differential cross section $d\tilde{\sigma}$ for the $2\to 3$ scattering processes, $\nu/\chi+\mathcal{N}\to\nu/\chi+\mathcal{N}+\gamma$, between $\nu/\chi$ and a stationary nucleus or nucleon $\mathcal{N}$ with mass $m_{\mathcal{N}}$. See Appendix A for calculational details.
@@
It is noteworthy that since the typical energy scale of the incoming $\nu/\chi$ that takes part in the $2\rightarrow 3$ process is $\mathcal{O}(0.1-1)$ GeV, and therefore, the energy and angular (cosine) spectra receive contributions from both coherent and incoherent regimes. The former regime is relevant for momentum transfers less than the de Broglie wavelength of the nucleus [$Q^{2}\lesssim(100\leavevmode\nobreak\ \text{MeV})^{2}$], whereas the latter corresponds to momentum transfers of the order of the proton radius. Here $Q^{2}=-(p_{\mathcal{N},{\rm out}}-p_{\mathcal{N},{\rm in}})^{2}$. The characteristic momenta range and the associated nucleus/nucleon responses for the coherent and incoherent regimes are imposed using the Helm form factor Lewin:1995rx ; Dobrich:2015jyk , $F_{\text{helm}}(Q^{2},A)$, and dipole form factor Kharzeev:2021qkd , $F_{\text{dipole}}(Q^{2})$, respectively.
@@
In the coherent regime, the Yukawa charge of the nucleus is treated as the coherent sum of the couplings to individual nucleons, whereas in the incoherent regime, nucleons contribute independently. Since the scalar couplings are defined in terms of fundamental quarks in Eq. (1), they can be related to nucleons using the form factors $f^{p,n}_{Tq}$ DeRomeri:2024iaw ; Cirelli:2013ufw , which are associated with the spin-independent operators defined in Ref. DelNobile:2021wmp . Therefore, the coherent and incoherent charges for a nucleus with an atomic number and mass $Z$ and $A$, where $N=A-Z$ is the number of neutrons, are,
@@
Figure 2: MiniBooNE and MicroBooNE fits for ${\rm neutrino}+{\rm dark}$ matter scattering where the mediators used are $m_{Z^{\prime}}=93.6$ MeV, and $m_{\phi}=87.6$ MeV, and the combination of couplings are $g=g_{\nu}g_{\phi Z^{\prime}\gamma}y_{q}=8.2\times 10^{-7}\leavevmode\nobreak\ \text{GeV}^{-1}$.
@@
Methodology. We begin our analysis by estimating the neutrino/dark matter flux at the BNB. For the neutrino fluxes at the detector, we adopt the published data in Ref. MiniBooNE:2008hfu for MiniBooNE, and Ref. MicroBooNE:2016pwy for MicroBooNE. As mentioned earlier, dark matter particles originate from the two-body decay of $Z^{\prime}$, which is produced predominantly via the neutrino leg in the three-body decay of charged mesons, ±±$\mathcal{M}^{\pm}\to\ell^{\pm}\nu_{l}Z^{\prime}$. Since the magnetic horn system focuses charged mesons before decaying, it is important to simulate the horn effects properly. To this end, “focused” charged-meson fluxes are generated using the RKHorn package RKHorn2024 . We then simulate their decays into $Z^{\prime}$ which subsequently decays into dark matter, collecting the resulting candidates directed toward the MiniBooNE and MicroBooNE detectors. Since the candidates reaching MiniBooNE and MicroBooNE are highly boosted along the beamline, we assume that their momentum is aligned to the beamline.
@@
We next perform Monte Carlo simulations to estimate the energy and angular spectra of the photon in the $2\to 3$ scattering process detailed in the previous section. The technical details for obtaining these kinematic distributions of final-state photons are described in the Appendix C. Since the cosine spectra of the MiniBooNE and MicroBooNE excesses have a striking forward and off-forward nature, we find that Yukawa couplings must equally accommodate momentum transfers in the coherent and incoherent regime. This can be ensured if the coherent and incoherent charges in Eq. (3) are somewhat at par. We accomplish this by invoking a mild cancelation ($\sim$ 10 – 30%) between $x_{u}$ and $x_{d}$ in the coherent charge that balances the $Z^{2}/N^{2}$ enhancement. For illustration, we choose an example of $x_{u}\sim 1$ and $x_{d}\sim-1.2$, so that the coherent and incoherent charges have similar magnitudes.
In our analysis, we compute the single-photon signal from neutrino and dark matter scattering separately and then combine them to fit with MiniBooNE. This depends on three mass parameters and four coupling parameters. However, since a produced $Z^{\prime}$ promptly decays into a $\chi$ pair for any sufficiently large $g_{\chi}$ (with our benchmark choice being $\alpha_{\chi}\equiv\frac{g_{\chi}^{2}}{4\pi}=0.5$), both experiments are insensitive to $g_{\chi}$. As a result, they are effectively sensitive to a single effective coupling, denoted by $g=g_{\nu}g_{\phi Z^{\prime}\gamma}y_{q}$. Adopting the conventional choice $m_{\chi}=m_{Z^{\prime}}/3$, we scan over the three parameters, $m_{Z^{\prime}}$, $m_{\phi}$, and $g$, to find the best-fit values (see Table 1) by minimizing $(\chi^{2}/\text{dof})_{\text{cosine}}+(\chi^{2}/\text{dof})_{\text{energy}}$ with the excess in MiniBooNE’s neutrino mode (see Appendix D for further analysis details).
${\rm Neutrino}+{\rm dark\leavevmode\nobreak\ matter\leavevmode\nobreak\ fits}$

### v2

@@ Abstract
Existing solutions to this anomaly mostly involve additional species in the neutrino sector, utilizing their oscillation, decays, and/or upscattering  Sorel:2003hf ; Karagiorgi:2009nb ; Collin:2016aqd ; Giunti:2011gz ; Giunti:2011cp ; Gariazzo:2017fdh ; Boser:2019rta ; Kopp:2011qd ; Kopp:2013vaa ; Dentler:2018sju ; Abazajian:2012ys ; Conrad:2012qt ; Diaz:2019fwt ; Asaadi:2017bhx ; Karagiorgi:2012kw ; Pas:2005rb ; Doring:2018cob ; Kostelecky:2003cr ; Katori:2006mz ; Diaz:2010ft ; Diaz:2011ia ; Gninenko:2009ks ; Gninenko:2009yf ; Bai:2015ztj ; Moss:2017pur ; Bertuzzo:2018itn ; Ballett:2018ynz ; Fischer:2019fbw ; Moulai:2019gpi ; Dentler:2019dhz ; deGouvea:2019qre ; Datta:2020auq ; Dutta:2020scq ; Abdallah:2020biq ; Abdullahi:2020nyr ; Liao:2016reh ; Carena:2017qhd ; Abdallah:2020vgg ; Abdallah:2024uby ; Hammad:2021mpl . Recently, long-lived mediators and dark matter upscattering have been considered to solve the MiniBooNE puzzle Dutta:2021cip ; PhysRevD.109.095017 . MicroBooNE’s strong capability to distinguish photons from electrons MicroBooNE:2025khi has already constrained the parameter space of some of these solutions, in particular, through a potential signature in the inclusive single-photon channel MicroBooNE:2025ntu but without any obvious excess in the coherent-photon production channel similar to SM MicroBooNE:2025rsd ; Hagaman:2025mrx . Already, these hints are beginning to inform a better understanding of the anomaly as it takes shape.
@@
Neutrinos and dark matter can arise from charged meson two- and three-body decays Dutta:2021cip , correlating their fluxes with the MiniBooNE target-mode excess via magnetic focusing in the BNB horn system. This mechanism simultaneously accounts for the absence of an excess in MiniBooNE’s dump-mode observations MiniBooNEDM:2018cxm , where the fluxes are isotropically suppressed due to the lack of magnetic focusing horns Dutta:2021cip .
@@
Model. We consider a simplified model which includes a massive scalar coupling to quarks, and a massive vector boson couples to either neutrinos, dark matter, or both simultaneously. Under this scenario, the scattering process delineated in Fig. 1 is expected to result in a single photon. A similar process occurs in the SM with neutrinos, via the $Z$-$h$-$\gamma$ interaction with the SM $Z$ and Higgs $h$. However, the associated cross section and kinematics from these heavy mediators are unsuitable to explain the MiniBooNE excess. This motivates us to introduce a new gauge boson, $Z^{\prime}$, and a scalar, $\phi$, lighter than the $Z$ and $h$. The new scalar $\phi$ can emerge from the extensions of the SM Higgs sector Dutta:2020scq ; Dutta:2022fdt , or as a new Higgs field in association with spontaneous breaking of the new $U(1)$ symmetry Dutta:2019fxn . The effective Lagrangian for $\phi$ and $Z^{\prime}$ with masses $m_{\phi}$ and $m_{Z^{\prime}}$, respectively, relevant for the $2\to 3$ neutrino scattering, is
$\mathcal{L}_{\nu}\supset g_{\nu}\bar{\nu}_{\ell}\gamma^{\alpha}\nu_{\ell}Z_{\alpha}^{\prime}-\frac{g_{\phi Z^{\prime}\gamma}}{2}\phi F_{\mu\nu}{Z^{\prime}}^{\mu\nu}+y_{\phi}\sum_{f}x_{f}\bar{f}f\phi,$ (1)
@@
where $\ell=e,\mu$, $F_{\mu\nu}$ ($Z^{\prime}_{\mu\nu}$) symbolizes the energy-stress tensor of the SM photon ($Z^{\prime}$) field, $g_{i}$ denotes the associated coupling, $y_{\phi}$ is the SM fermion flavor-universal coupling, and $x_{f}$ parameterizes the flavor-nonuniveral Yukawa fractions. For the $\nu$-$\phi$ interaction, one may consider either the Dirac type $\bar{\nu}_{\ell}\nu_{R}\phi$ or the Majorana type $\bar{\nu}_{\ell}^{c}\nu_{\ell}\phi$. For the light dark matter solution, we use an alternative (but similar) Lagrangian that generates production of dark natter via $Z^{\prime}$ and/or $\phi$, as well as $\chi\mathcal{N}\to\chi\mathcal{N}\gamma$ for a fermion $\chi$ with mass $m_{\chi}$:
$\mathcal{L}_{\chi}\supset g_{\chi}\bar{\chi}\gamma^{\mu}\chi Z^{\prime}_{\mu}+y_{\chi}\bar{\chi}\chi\phi-\frac{g_{\phi Z^{\prime}\gamma}}{2}\phi F_{\mu\nu}{Z^{\prime}}^{\mu\nu}+y_{\phi}\sum_{f}x_{f}\bar{f}f\phi$.
@@
In this scenario, $\chi$ is predominantly produced through exotic three-body decays of charged mesons to explain the anomaly, followed by its decays into $\chi$ pairs. Example UV-complete models realizing the above interactions will be discussed later in the paper. Based on the above Lagrangians, we calculate the differential cross section $d\tilde{\sigma}$ for the $2\to 3$ scattering processes, $\nu/\chi+\mathcal{N}\to\nu/\chi+\mathcal{N}+\gamma$, between $\nu/\chi$ and a stationary nucleus or nucleon $\mathcal{N}$ with mass $m_{\mathcal{N}}$. See Appendix A for calculational details.
@@
It is noteworthy that since the typical energy scale of the incoming $\nu/\chi$ that takes part in the $2\rightarrow 3$ process is $\mathcal{O}(0.1-1)$ GeV, and therefore, the energy and angular (cosine) spectra receive contributions from both coherent and incoherent regimes. The former regime is relevant for momentum transfers less than the de Broglie wavelength of the nucleus $Q^{2}\lesssim(100\leavevmode\nobreak\ \text{MeV})^{2}$], whereas the latter corresponds to momentum transfers of the order of the proton radius. Here $Q^{2}=-(p_{\mathcal{N},{\rm out}}-p_{\mathcal{N},{\rm in}})^{2}$. The characteristic momenta range and the associated nucleus/nucleon responses for the coherent and incoherent regimes are imposed using the Helm form factor Lewin:1995rx ; Dobrich:2015jyk , $F_{\text{helm}}(Q^{2},A)$, and dipole form factor Kharzeev:2021qkd , $F_{\text{dipole}}(Q^{2})$, respectively.
@@
In the coherent regime, the Yukawa charge of the nucleus is treated as the coherent sum of the couplings to individual nucleons, whereas in the incoherent regime, nucleons contribute independently. Since the scalar couplings are defined in terms of fundamental quarks in Eq. (1), they can be related to nucleons using the form factors $f^{p,n}_{Tq}$ DeRomeri:2024iaw ; Cirelli:2013ufw , which are associated with the spin-independent operators defined in Ref. DelNobile:2021wmp . Therefore, the coherent and incoherent charges for a nucleus with an atomic number and mass $Z$ and $A$, where N=AZN=A-Z is the number of neutrons, are,
@@
Figure 2: MiniBooNE fits (left) and MicroBooNE overlays (right) with dark matter coupled to a $Z^{\prime}$ and $\phi$ exchanged with the nucleus where $m_{Z^{\prime}}=93.6$ MeV, $m_{\phi}=87.6$ MeV, $m_{\chi}=m_{Z^{\prime}}/3$ and $g_{D}=0.5$. After maximizing the dark matter flux from meson decays, the fits require $g=g_{\phi Z^{\prime}\gamma}y_{\phi}=5\times 10^{-4}\leavevmode\nobreak\ \text{GeV}^{-1}$. The black and colored error bars depict the statistical and assumed systematic errors, respectively.
@@
Methodology. We begin our analysis by estimating the neutrino/dark matter flux at the BNB. For the neutrino fluxes at the detector, we adopt the published data in Ref. MiniBooNE:2008hfu for MiniBooNE, and Ref. MicroBooNE:2016pwy for MicroBooNE. Dark matter particles originate from mediators, which originate from charged mesons, such as $\mathcal{M}^{\pm}\to\ell^{\pm}+\nu_{l}+Z^{\prime}/\phi$. Since the magnetic horn system focuses charged mesons before they decay, it is essential to simulate the horn effects accurately. To this end, “focused” charged-meson fluxes are generated using the RKHorn package RKHorn2024 . We then simulate their decays into $Z^{\prime}/\phi$, which subsequently decay into dark matter, collecting the resulting candidates directed toward the MiniBooNE and MicroBooNE detectors. Since the candidates reaching MiniBooNE and MicroBooNE are highly boosted along the beamline, we assume that they are aligned with the beamline. We find that the energy flux of the dark matter reaching the detector is largely independent of the Lorentz structure of the mediator, and the mean energy is generally $\sim$300-400 MeV less than that of the neutrino flux.
@@
We next perform Monte Carlo simulations to estimate the energy and angular spectra of the photon in the $2\to 3$ scattering process detailed in the previous section. The technical details for obtaining these kinematic distributions of final-state photons are described in the Appendix C. Since the cosine spectra of the MiniBooNE and MicroBooNE excesses have a striking forward and off-forward nature, we find that Yukawa couplings must equally accommodate momentum transfers in the coherent and incoherent regimes. This can be ensured if the coherent and incoherent charges in Eq. (3) are somewhat at par. We accomplish this by invoking a mild cancelation ($\sim$ 10 – 30%) between $x_{u}$ and $x_{d}$ in the coherent charge that balances the $Z^{2}\leavevmode\nobreak\ \rm{or}\leavevmode\nobreak\ N^{2}$ enhancement. For illustration, we choose an example of $x_{u}\sim 1$ and $x_{d}\sim-1.2$, so that the coherent and incoherent charges have similar magnitudes.
In our analysis, we compute the single-photon signal from neutrino and dark matter scattering separately. The neutrino fits depend on two mass parameters ($m_{\phi}$ and $m_{Z^{\prime}}$) and three couplings. However, they can be effectively reduced to three, as the fit is sensitive to the product of couplings, i.e., $g\equiv g_{\nu}g_{\phi Z^{\prime}\gamma}y_{\phi}$. In contrast, the dark matter fits additionally depend on $m_{\chi}$ and $g_{\chi}$. The specific choices of these additional parameters do not affect our main conclusion, although we do keep $m_{\chi}$ light enough as to not slow down the dark matter propagation outside of the MiniBooNE timing constraint MiniBooNE:2020pnu . For illustration, we adopt the conventional choice $m_{\chi}=\dfrac{m_{Z^{\prime}}}{3}$ and $\alpha_{\chi}\equiv\dfrac{g_{\chi}^{2}}{4\pi}=0.5$. Assuming that the mediator production from charged-meson decays is maximized—thereby maximizing the dark matter flux—the dark matter fits depend on three parameters, i.e., $m_{\phi},m_{Z^{\prime}},$ and $g\equiv g_{\phi Z^{\prime}\gamma}y_{\phi}$. Since more data are available from MiniBooNE’s neutrino mode, we scan over the three parameters, $m_{Z^{\prime}}$, $m_{\phi}$, and $g$, to find the best-fit values (see Table 1) by minimizing $(\chi^{2}/\text{dof})_{\text{cosine}}+(\chi^{2}/\text{dof})_{\text{energy}}$ with the excess in MiniBooNE’s neutrino mode (see Appendix. D). Using MiniBooNE neutrino mode’s best-fit parameters, we find the overlay of the new physics signal with MiniBooNE’s antineutrino mode and MicroBooNE neutrino mode excess.
Dark matter fits ($\alpha_{\chi}=0.5$)
$g_{\phi Z^{\prime}\gamma}y_{\phi}$ Average $\chi^{2}/$dof
[…]
1 (45.41, 48.49, 16.17) $\times 10^{-4} 2.2$0 1.28
2 (87.67, 93.63, 31.21) $5\times 10^{-4}$ 1.89 1.27
3 (119.08, 114.06, 38.02) $7\times 10^{-4}$ 2.21 1.34
4 (193.07, 138.95, 46.31) $1\times 10^{-3}$ 1.79 1.41
$\rm Neutrino\leavevmode\nobreak\ fits$}